OpenStudy (anonymous):

Which function below has the lowest y-intercept? f(x) graph going through (0,5) and (4, 3) g(x) Kyle started the summer having read 3 books but plans to read 6 books per month over the summer. h(x) h(x) = 3x + 4 f(x) g(x) h(x) The functions all have the same y-intercept.

OpenStudy (anonymous):

Help please :c

OpenStudy (ybarrap):

We know that linear equations have a slope (m) and y-intercept (b): y=mx+b, where x and y are variables. One point for f(x) is <0,5>, where x=0 and y=5. Another point is <4,3>. So we have two equations and two unknowns. We then can solve for m and b: $$ 5=m(0)+b\\ 4=m(4)+b $$ Using the 1st equation, we immediately see that b=5. This is the y=intercept for f(x). Make sense?

OpenStudy (anonymous):

Okay, so then I would do 6 = m(3) + b for g(x)?

OpenStudy (ybarrap):

The second equation should have been $$ 4=m(3)+b $$

OpenStudy (ybarrap):

Good guess. But saying that she "started" the summer reading 3 books, means that at x=0, y=3, using our equation notation: 3=m(0) + b. This immediately shows the y-intercept for g(x).

OpenStudy (anonymous):

b = 3 ?

OpenStudy (anonymous):

for g(x)

OpenStudy (ybarrap):

yes

OpenStudy (anonymous):

Okay, just one more question. How would I find h(x)?

OpenStudy (ybarrap):

So you are given h(x): y=3x+5. Here m=3 and b=5. So you are essentially just given the y-intercept, b. Collect all these y-intercepts and find the lowest one. y-intercepts: f(x) has b=? g(x) has b=? h(x) has b=?

OpenStudy (anonymous):

So g(x) should be the answer correct?

OpenStudy (ybarrap):

note that h(x) just means " A function, let's call it 'h' and it is a function of 'x' " So, h(x)=y=3x+4.

OpenStudy (ybarrap):

Why?

OpenStudy (anonymous):

I'm doing this for homework and just confused by this all

OpenStudy (ybarrap):

do you understand what we did here?

OpenStudy (anonymous):

Yeah I do, the graph was obvious to me but the other two I was confused about

OpenStudy (ybarrap):

that's ok. Sometimes you need to see the problem in different ways to see that it really is all the same thing. In this case, everything boils down to y=mx+b, which is just a graph, with slope=m and that crosses the y-axis at y=b, when x=0.

OpenStudy (anonymous):

Okay, I think I got it, thanks for your time! :)

OpenStudy (ybarrap):

No problem. Hope you start to enjoy math, it can be really fun.